### v1

# Topology-aware Embedding Memory for Learning on Expanding Graphs

Abstract.
[…]
copyright: nonecopyright: Nonejournalyear: 2018doi: XXXXXXX.XXXXXXX
1. Introduction
Traditional graph learning works typically assume the graph to be static. However, real-world graphs often expand constantly with emerging new types of nodes and their associated edges. Accordingly, models trained incrementally on the new node types may experience catastrophic forgetting (severe performance degradation) on the old ones. Targeting this challenge, continual learning on expanding graphs (Liu et al., 2021; Zhou and Cao, 2021; Zhang et al., 2022b) attracts increasingly more attention recently. It exhibits enormous value in various practical applications, especially in the case where graphs are relatively large, and retraining a new model over the entire graph is computationally infeasible. For instance, in a social network, a community detection model has to keep adapting its parameters based on nodes from newly emerged communities; in a citation network, a document classifier needs to continuously update its parameters to distinguish the documents of newly emerged research fields.
Memory replay (Rebuffi et al., 2017; Lopez-Paz and Ranzato, 2017; Aljundi et al., 2019; Shin et al., 2017), which stores representative samples in a buffer to retrain the model and maintain its performance over existing tasks, exhibits great success in preventing catastrophic forgetting for various continual learning tasks, e.g., computer vision and reinforcement learning (Kirkpatrick et al., 2017; Li and Hoiem, 2017; Aljundi et al., 2018; Rusu et al., 2016). Directly applying memory replay to graph data with the popular message passing neural networks (MPNNs, the general framework for most GNNs) (Gilmer et al., 2017; Kipf and Welling, 2017; Veličković et al., 2017), however, could give rise to the memory explosion problem because the necessity to consider the explicit topological information of target nodes. Specifically, due to the message passing over the topological connections in graphs, retraining an $L$-layer GNN (Figure 1, left) with $n$ buffered nodes would require storing $\mathcal{O}(nd^{L})$ nodes (Chiang et al., 2019; Chen et al., 2017) (the number of edges is not counted yet) in the buffer, where $d$ is the average node degree.
Figure 1. (a) ER-GNN (Zhou and Cao, 2021) that stores the input attributes of individual nodes. (b) Sparsified Subgraph Memory (SSM) (Zhang et al., 2022c) that stores sparsified computation ego-subgraphs. (c) Our PDGNNs with TEM. The incoming computation ego-subgraphs are embedded as TEs and then fed into the trainable function. The stored TEs are sampled based on their coverage ratio (Section 3.6).
Take the Reddit dataset (Hamilton et al., 2017) as an example, its average node degree is 492, and the buffer size will easily be intractable even with a 2-layer GNN. To resolve this issue, Experience Replay based GNN (ER-GNN) (Zhou and Cao, 2021) stores representative input nodes (i.e., node attributes) in the buffer but completely ignores the topological information (Figure 1 a). Feature graph network (FGN) (Wang et al., 2020a) implicitly encodes node proximity with the inner products between the features of the target node and its neighbors. However, the explicit topological connections are abandoned and message passing is no longer feasible on the graph. Sparsified Subgraph Memory (SSM) (Zhang et al., 2022c) sparsifies the computation ego-subgraphs for tractable memory consumption, which still partially sacrifices topological information, especially when the computation ego-subgraphs are large and a majority of nodes/edges is removed after sparsification (Figure 1 b).
To this end, we present a general framework of Parameter Decoupled Graph Neural Networks (PDGNNs) with Topology-aware Embedding Memory (TEM) to perform continual learning on expanding graphs (Figure 1 c). First, with systematic analysis, we demonstrate that the necessity to store the complete computation ego-subgraphs for retraining MPNNs arises from the entanglement between the trainable parameters and the individual nodes/edges (Section 3.2). Targeting this problem, we design the Parameter Decoupled Graph Neural Networks (PDGNNs) framework, which decouples the trainable parameters from individual nodes/edges. The PDGNNs framework enables us to develop a novel concept, Topology-aware Embedding (TE), which is a vector with a fixed size but contains all necessary information for retraining the trainable parameters of PDGNNs. Such TEs are desired surrogates of computation ego-subgraphs to facilitate memory replay. After learning each task, a subset of TEs is selected with a certain sampling strategy and stored in the Topology-aware Embedding Memory (TEM). Because the size of a TE is fixed, the memory space complexity of a buffer (with size $n$) can be dramatically reduced from $\mathcal{O}(nd^{L})$ to $\mathcal{O}(n)$. Moreover, different from traditional continual learning on Euclidean data without explicit topological connections (e.g., images), we theoretically discover that replaying the TE of one single node incurs a pseudo-training effect on the neighboring nodes, which could also alleviate the forgetting problem for the other nodes within the same computation ego-subgraph. This unique phenomenon in continual learning on expanding graphs takes place due to the neighborhood aggregation in GNNs. The pseudo-training effect suggests that TEs corresponding to larger computation ego-subgraphs (quantitatively measured by coverage ratio) are more beneficial to continual learning performance. Based on the theoretical finding, we develop a novel coverage maximization sampling strategy, which enlarges the coverage ratio of the selected TEs and empirically enhances the performance for a tight memory budget. In our experiments, we adopt both the class-incremental (class-IL) continual learning scenario (Rebuffi et al., 2017; Zhang et al., 2022c, a) and the task-incremental (task-IL) scenario (Liu et al., 2021; Zhou and Cao, 2021), to evaluate the effectiveness of PDGNNs-TEM. Thorough empirical studies demonstrate that PDGNNs with TEM outperform the state-of-the-art techniques in both class-IL and task-IL scenarios.
2. Related Works
2.1. Continual Learning & Continual Learning on Expanding Graphs
To alleviate the catastrophic forgetting problem, existing approaches can be categorized into regularization, memory replay, and parameter isolation based methods. Regularization based methods aim to prevent drastic modification to parameters that are important for previous tasks (Farajtabar et al., 2020; Kirkpatrick et al., 2017; Li and Hoiem, 2017; Aljundi et al., 2018; Hayes and Kanan, 2020; Rakaraddi et al., 2022; Sun et al., 2022; Qi et al., [n. d.]). Parameter isolation methods adaptively allocate new parameters for the new tasks to protect the ones for the previous tasks (Wortsman et al., 2020; Wu et al., 2019a; Yoon et al., 2020, 2017; Rusu et al., 2016). Memory replay based methods store and replay representative data from previous tasks when learning new tasks (Caccia et al., 2020; Chrysakis and Moens, 2020; Rebuffi et al., 2017; Lopez-Paz and Ranzato, 2017; Aljundi et al., 2019; Shin et al., 2017). Recently, continual learning on expanding graphs attracts increasingly more attention due to its practical importance  (Das and Isufi, 2022a, c, b; Zhou and Cao, 2021; Liu et al., 2021; Wang et al., 2020b; Xu et al., 2020; Daruna et al., 2021; Carta et al., 2021; Zhang et al., 2022c, b, a; Febrinanto et al., 2023; Yuan et al., 2023; Cai et al., 2022; Wei et al., 2022; Wang et al., 2022; Ahrabian et al., 2021; Su and Wu, 2023; Kim et al., 2022; Daruna et al., 2021; Hoffmann et al., 2023; Galke et al., 2023; Li et al., 2023; Cui et al., 2023). Existing methods include regularization ones like topology-aware weight preserving (TWP) (Liu et al., 2021) that preserves crucial topologies, parameter isolation methods like HPNs (Zhang et al., 2022b) that select different parameters for different tasks, and memory replay methods like ER-GNN (Zhou and Cao, 2021) and Sparsified Subgraph Memory (SSM) (Zhang et al., 2022c) that store representative nodes or sparsified computation ego-subgraphs. Our work is also memory based and its key advantage is the capability to preserve complete topological information with reduced space complexity, which shows significant superiority in class-IL setting (Section 4.4). Finally, it is worth highlighting the difference between continual learning on expanding graphs and some relevant research areas. First, dynamic graph learning (Galke et al., 2021; Wang et al., 2020d; Han et al., 2020; Yu et al., 2018; Nguyen et al., 2018; Zhou et al., 2018; Ma et al., 2020; Feng et al., 2020) focuses on the temporal dynamics with all previous data being accessible. In contrast, continual learning on expanding graphs aims to alleviate forgetting, therefore the previous data is inaccessible. Second, few-shot graph learning (Zhou et al., 2019; Guo et al., 2021; Yao et al., 2020; Tan et al., 2022) targets fast adaptation to new tasks. In training, few-shot learning models can access all previous tasks (unavailable in continual learning). In testing, few-shot learning models need to be fine-tuned on the test classes, while the continual learning models are tested on existing tasks without fine-tuning.
2.2. Graph Neural Networks &\&& Reservoir Computing
Graph Neural Networks (GNNs) are deep learning models designed to generate representations for graph data, which typically interleave the neighborhood aggregation and node feature transformation to extract the topological features (Kipf and Welling, 2017; Gilmer et al., 2017; Veličković et al., 2017; Xu et al., 2018; Chen et al., 2018; Hamilton et al., 2017; Yan et al., 2022; Yang et al., 2023; Zhang et al., 2022d). GNNs without interleaving the neighborhood aggregation and node feature transformation have been developed to reduce the computation complexity and increase the scalability (Zeng et al., 2021; Chen et al., 2020, 2019; Nt and Maehara, 2019; Frasca et al., 2020; Fey et al., 2021; Cong et al., 2020; Dong et al., 2021). For example, Simple Graph Convolution (SGC) (Wu et al., 2019b) removes the non-linear activation from GCN (Kipf and Welling, 2017) and only keeps one neighborhood aggregation and one node transformation layer. Approximate Personalized Propagation of Neural Predictions (APPNP) (Klicpera et al., 2018) first performs node transformation and then conducts multiple neighborhood aggregations in one layer. Motivated by these works, the PDGNNs framework in this paper is specially designed to decouple the neighborhood aggregation with trainable parameters, and derive the topology-aware embeddings (TEs) to reduce the memory space complexity and facilitate continual learning on expanding graphs. Besides, PDGNNs are also related to reservoir computing (Gallicchio and Micheli, 2020, 2010), which can embed the input data (e.g. graphs) via a fixed non-linear system. The reservoir computing modules can be adopted in PDGNNs as the TE generation function (Equation 4), and the corresponding experiments are included in Appendix B.1.
3. Parameter Decoupled GNNs with Topology-aware Embedding Memory
In this section, we first introduce the notations, and then explain the technical challenge of applying memory replay techniques to GNNs. Targeting the challenge, we introduce PDGNNs with Topology-aware Embedding Memory (TEM). Finally, inspired by theoretical findings of the pseduo-training effect, we develop the coverage maximization sampling to enhance the performance when the memory budget is tight, which has shown its effectiveness in our empirical study. All detailed proofs are provided in Appendix A.
3.1. Preliminaries
Continual learning on expanding graphs is formulated as learning node representations on a sequence of subgraphs (tasks): $\mathcal{S}$ = $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{T}\}$. Each $\mathcal{G}_{\tau}$ (i.e., $\tau$-th task) contains several new categories of nodes in the overall graph, and is associated with a node set $\mathbb{V}_{\tau}$ and an edge set $\mathbb{E}_{\tau}$, which is represented as the adjacency matrix $\mathbf{A}_{\tau}\in\mathbb{R}^{|\mathbb{V}_{\tau}|\times|\mathbb{V}_{\tau}|}$. The degree of a node $d$ refers to the number of edges connected to it. In practice, $\mathbf{A}_{\tau}$ is often normalized as $\hat{\mathbf{A}}_{\tau}=\mathbf{D}_{\tau}^{-\frac{1}{2}}\mathbf{A}_{\tau}\mathbf{D}_{\tau}^{-\frac{1}{2}}$, where $\mathbf{D}_{\tau}\in\mathbb{R}^{|\mathbb{V}_{\tau}|\times|\mathbb{V}_{\tau}|}$ is the degree matrix. Each node $v\in\mathbb{V}_{\tau}$ has a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{b}$. In classification tasks, each node $v$ has a label $\mathbf{y}_{v}\in\{0,1\}^{C}$, where $C$ is the total number of classes. When generating the representation for a target node $v$, a L-layer GNN typically takes a computation ego-subgraph $\mathcal{G}^{sub}_{{\tau},v}$, containing the $L$-hop neighbors of $v$ (i.e. $\mathcal{N}^{L}(v)$), as the input. For simplicity, $\mathcal{G}^{sub}_{v}$ is used in the following, without the graph index.
3.2. Memory Replay Meets GNNs
In traditional continual learning, a model $\mathrm{f}(\cdot;\bm{\theta})$ parameterized by $\bm{\theta}$ is sequentially trained on $T$ tasks. Each task $\tau$ ($\tau\in\{1,...,T\}$) corresponds to a dataset $\mathbb{D}_{\tau}=\{(\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n_{\tau}}\}$. To avoid forgetting, memory replay based methods store representative data from the old tasks in a buffer $\mathcal{B}$. When learning new tasks. A common approach to utilize $\mathcal{B}$ is through an auxiliary loss:
(1) $\mathcal{L}=\underbrace{\sum_{\mathbf{x}_{i}\in\mathbb{D}_{\tau}}l(\mathrm{f}(\mathbf{x}_{i};\bm{\theta}),\mathbf{y}_{i})}_{\text{$\mathcal{L}_{\tau}$: loss of the current task}}+\lambda\underbrace{\sum_{\mathbf{x}_{j}\in\mathcal{B}}l(\mathrm{f}(\mathbf{x}_{j};\bm{\theta}),\mathbf{y}_{j})}_{\text{$\mathcal{L}_{aux}$: auxiliary loss}},$
where $l(\cdot,\cdot)$ denotes the loss function, and $\lambda\geq 0$ balances the contribution of the old data. Instead of directly minimizing $\mathcal{L}_{aux}$, the buffer $\mathcal{B}$ may also be used in other ways to prevent forgetting (Lopez-Paz and Ranzato, 2017; Rebuffi et al., 2017). In these applications, the space complexity of a buffer containing $n$ examples is $\mathcal{O}(n)$.
However, to capture the topological information, GNNs obtain the representation of a node $v$ based on a computation ego-subgraph surrounding $v$. We exemplify it with the popular MPNN framework (Gilmer et al., 2017), which updates the hidden node representations at the $l+1$-th layer as:
(2) $\mathbf{m}_{v}^{l+1}=\sum_{w\in\mathcal{N}^{1}(v)}\mathrm{M}_{l}(\mathbf{h}^{l}_{v},\mathbf{h}^{l}_{w},\mathbf{x}^{e}_{v,w};\bm{\theta}_{l}^{\mathrm{M}}),$ $\mathbf{h}^{l+1}_{v}=\mathrm{U}_{l}(\mathbf{h}_{v}^{l},\mathbf{m}_{v}^{l+1};\bm{\theta}_{l}^{\mathrm{U}}),$
where $\mathbf{h}_{v}^{l}$, $\mathbf{h}_{w}^{l}$ are hidden representations of nodes at layer $l$, $\mathbf{x}^{e}_{v,w}$ is the edge feature, $\mathrm{M}_{l}(\cdot,\cdot,\cdot;\bm{\theta}_{l}^{\mathrm{M}})$ is the message function to integrate neighborhood information, and $\mathrm{U}_{l}(\cdot,\cdot;\bm{\theta}_{l}^{\mathrm{U}})$ updates $\mathbf{m}_{v}^{l+1}$ into $\mathbf{h}_{v}^{l}$ ($\mathbf{h}_{v}^{0}$ is the input features). In a $L$-layer MPNN, the representation of a node $v$ can be simplified as,
(3) $\mathbf{h}^{L}_{v}=\mathrm{MPNN}(\mathbf{x}_{v},\mathcal{G}^{sub}_{v};\bm{\Theta}),$
where $\mathcal{G}^{sub}_{v}$ contains the $L$-hop neighbors ($\mathcal{N}^{L}(v)$), $\mathrm{MPNN}(\cdot,\cdot;\bm{\Theta})$ is the composition of all $\mathrm{M}_{l}(\cdot,\cdot,\cdot;\bm{\theta}_{l}^{\mathrm{M}})$ and $\mathrm{U}_{l}(\cdot,\cdot;\bm{\theta}_{l}^{\mathrm{U}})$ at different layers. Since $\mathcal{N}^{L}(v)$ typically contains $\mathcal{O}(d^{L})$ nodes, replaying $n$ nodes requires storing $\mathcal{O}(nd^{L})$ nodes (the edges are not counted yet), where $d$ is the average degree. Therefore, the buffer size will be easily intractable in practice (e.g. the example of Reddit dataset in Introduction), and directly storing the computation ego-subgraphs for memory replay is infeasible for GNNs.
3.3. Parameter Decoupled GNNs with TEM
As we discussed earlier, the key challenge of applying memory replay to graph data is to preserve the rich topological information of the computation ego-subgraphs with potentially unbounded sizes. Therefore, a natural resolution is to preserve the crucial topological information with a compact vector such that the memory consumption is tractable. Formally, the desired subgraph representation can be defined as Topology-aware Embedding (TE).
Definition 0 (Topology-aware embedding).
Given a specific GNN parameterized with $\bm{\theta}$ and an input $\mathcal{G}^{sub}_{v}$, an embedding vector $\mathbf{e}_{v}$ is a topology-aware embedding for $\mathcal{G}^{sub}_{v}$ with respect to this GNN, if optimizing $\bm{\theta}$ with $\mathcal{G}^{sub}_{v}$ or $\mathbf{e}_{v}$ for this specific GNN are equivalent, i.e. $\mathbf{e}_{v}$ contains all necessary topological information of $\mathcal{G}^{sub}_{v}$ for training this GNN.
However, TEs cannot be directly derived from the MPNNs due to their interleaved neighborhood aggregation and feature transformations. According to Section 3.2, whenever the trainable parameters get updated, recalculating the representation of a node $v$ requires all nodes and edges in $\mathcal{G}^{sub}_{v}$. To resolve this issue, we formulate the Parameter Decoupled Graph Neural Networks (PDGNNs) framework, which decouples the trainable parameters from the individual nodes/edges. PDGNNs may not be the only feasible framework to derive TEs, but is the first attempt and is empirically effective. Given $\mathcal{G}^{sub}_{v}$, the prediction of node $v$ with PDGNNs consists of two steps. First, the topological information of $\mathcal{G}^{sub}_{v}$ is encoded into an embedding $\mathbf{e}_{v}$ via the function $\mathrm{f}_{topo}(\cdot)$ without trainable parameters (instantiations of $\mathrm{f}_{topo}(\cdot)$ are detailed in Section 3.4).
(4) $\mathbf{e}_{v}=\mathrm{f}_{topo}(\mathcal{G}^{sub}_{v}).$
Next, $\mathbf{e}_{v}$ is further passed into a trainable function $\mathrm{f}_{out}(\cdot;\bm{\theta})$ parameterized by $\bm{\theta}$ (instantiations of $\mathrm{f}_{out}(\cdot;\bm{\theta})$ are detailed in Section 3.4) to get the output prediction $\mathbf{\hat{y}}_{v}$,
(5) $\mathbf{\hat{y}}_{v}=\mathrm{f}_{out}(\mathbf{e}_{v};\bm{\theta}).$
With the formulations above, $\mathbf{e}_{v}$ derived in Eq. (4) clearly satisfies the requirements of TE (Definition 3.1). Specifically, since the trainable parameters acts on $\mathbf{e}_{v}$ instead of any individual node/edge, optimizing the model parameters $\bm{\theta}$ with either $\mathbf{e}_{v}$ or $\mathcal{G}^{sub}_{v}$ are equivalent. Therefore, to retrain the model, the memory buffer only needs to store TEs instead of the original computation ego-subgraphs, which reduces the space complexity from $\mathcal{O}(nd^{L})$ to $\mathcal{O}(n)$. We name the buffer to store the TEs as Topology-aware Embedding Memory ($\mathcal{TEM}$). Given a new task $\tau$, the update of $\mathcal{TEM}$ is:
(6) $\mathcal{TEM}=\mathcal{TEM}\;\bigcup\;\mathrm{sampler}(\{\mathbf{e}_{v}\mid v\in\mathbb{V}_{\tau}\},n),$
where $\mathrm{sampler}(\cdot,\cdot)$ is the adopted sampling strategy to populate the buffer, $\bigcup$ denotes the set union, and $n$ is the budget. According to the experimental results (Section 4.3), as long as $\mathcal{TEM}$ is maintained, PDGNNs-TEM can perform reasonably well with different choices of $\mathrm{sampler}(\cdot,\cdot)$, including the random sampling. Nevertheless, in Section 3.6, based on the theoretical insights in Section 3.5, we propose a novel sampling strategy to better populate $\mathcal{TEM}$ when the memory budget is tight, which is empirically verified to be effective in Section 4.3. Besides, Equation (6) assumes that all data of the current task are presented concurrently. In practice, the data of a task may come in multiple batches (e.g., nodes come in batches on large graphs), and the buffer update have to be slightly modified, as detailed in Appendix C.1. For task $\tau$ with graph $\mathcal{G}_{\tau}$, the loss with $\mathcal{TEM}$ then becomes:
(7) $\mathcal{L}=\underbrace{\sum_{v\in\mathbb{V}_{\tau}}l(\mathrm{f}_{out}(\mathbf{e}_{v};\bm{\theta}),\mathbf{y}_{v})}_{\text{$\mathcal{L}_{\tau}$: loss of the current task $\tau$}}+\underbrace{\lambda\sum_{\mathbf{e}_{w}\in\mathcal{TEM}}l(\mathrm{f}_{out}(\mathbf{e}_{w};\bm{\theta}),\mathbf{y}_{w})}_{\text{$\mathcal{L}_{aux}$: auxiliary loss}}.$
$\lambda$ balances the contribution of the data from the current task and the memory, and is typically manually chosen in traditional continual learning works. However, on graph data, we adopt a different strategy to re-scale the losses according to the class sizes to counter the bias from the severe class imbalance, which cannot be handled on graphs by directly balancing the datasets (details are provided in Appendix B.6).
3.4. Instantiations of PDGNNs
Although without trainable parameters, the function $\mathrm{f}_{topo}(\cdot)$ for generating TEs can be highly expressive with various formulations including linear and non-linear ones, both of which are studied in this work. First, the linear instantiations of $\mathrm{f}_{topo}(\cdot)$ can be generally formulated as,
(8) $\mathbf{e}_{v}=\mathrm{f}_{topo}(\mathcal{G}^{sub}_{v})=\sum_{w\in\mathbb{V}}\mathbf{x}_{w}\cdot\pi(v,w;\hat{\mathbf{A}}),$
where $\pi(\cdot,\cdot;\hat{\mathbf{A}})$ denotes the strategy for computation ego-subgraph construction and determines how would the model capture the topological information. Equation 8 describes the operation on each node. In practice, Equation 8 could be implemented as matrix multiplication to generate TEs of a set of nodes $\mathbb{V}$ in parallel, i.e. $\mathbf{E}_{\mathbb{V}}=\bm{\Pi}\mathbf{X}_{\mathbb{V}}$, where each entry $\bm{\Pi}_{v,w}=\pi(v,w;\hat{\mathbf{A}})$. $\mathbf{E}_{\mathbb{V}}\in\mathbb{R}^{|\mathbb{V}|\times b}$ is the concatenation of all TEs ($\mathbf{e}_{v}\in\mathbb{R}^{b}$), and $\mathbf{X}_{\mathbb{V}}\in\mathbb{R}^{|\mathbb{V}|\times b}$ is the concatenation of all node feature vectors $\mathbf{x}_{v}\in\mathbb{R}^{b}$. In our experiments, we adopt three representative strategies. The first strategy (S1) (Wu et al., 2019b) is a basic version of message passing and can be formulated as $\bm{\Pi}=\hat{\mathbf{A}}^{L}$. The second strategy (S2) (Zhu and Koniusz, 2020) considers balancing the contribution of neighborhood information from different hops via a hyperparameter $\alpha$, i.e. $\bm{\Pi}=\frac{1}{L}\sum_{l=1}^{L}\big{(}(1-\alpha)\hat{\mathbf{A}}^{l}+\alpha\mathbf{I}\big{)}$. Finally, we also adopt a strategy (S3) (Klicpera et al., 2018) that adjusts the contribution of the neighbors based on PageRank (Page et al., 1999), i.e. $\bm{\Pi}=\big{(}(1-\alpha)\hat{\mathbf{A}}+\alpha\mathbf{I}\big{)}^{L}$, in which $\alpha$ also balances the contribution of the neighborhood information.
The linear formulation of $\mathrm{f}_{topo}(\cdot)$ (Equation (8)) yields both promising experimental results (Section B) and instructive theoretical results (Section 3.5, and 3.6). Equation (8) is also highly efficient especially for large graphs due to the absence of iterative neighborhood aggregations. But $\mathrm{f}_{topo}(\cdot)$ can also take non-linear forms with more complex mappings, e.g., the reservoir computing modules (Gallicchio and Micheli, 2020). The corresponding experimental and theoretical effects are introduced in Appendix B.1 and A.3.
Since $\mathrm{f}_{out}(\cdot;\bm{\theta})$ simply deals with individual vectors (TEs), it is instantiated as MLP in this work. The specific configurations of $\mathrm{f}_{out}(\cdot;\bm{\theta})$ is described in the experimental part (Section 4.2).
Figure 2. Illustration of the coverage ratio. Supposing the graph has $N$ nodes, $R_{c}(\{u\})=\frac{13}{N}$, $R_{c}(\{v\})=\frac{15}{N}$, $R_{c}(\{u\})=\frac{14}{N}$, and $R_{c}(\{u,v,w\})=\frac{42}{N}$
3.5. Pseudo-training Effects of TEs
In traditional continual learning on Euclidean data without explicit topological connections, replaying an example $\mathbf{x}_{i}$ (e.g., an image) only reinforces the prediction of $\mathbf{x}_{i}$ itself. In this subsection, we introduce the pseudo-training effect, which implies that training PDGNNs with $\mathbf{e}_{v}$ of node $v$ also influences the predictions of the other nodes in $\mathcal{G}^{sub}_{v}$, based on which we develop a novel sampling strategy to further boost the performance with a tight memory budget.
Theorem 1 (Pseudo-training).
Given a node $v$, its computation ego-subgraph $\mathcal{G}^{sub}_{v}$, the TE $\mathbf{e}_{v}$, and label $\mathbf{y}_{v}$ (suppose $v$ belongs to class $k$, i.e. $\mathbf{y}_{v,k}=1$), then training PDGNNs with $\mathbf{e}_{v}$ has the following two properties:
[…]
The pseudo-training effect essentially arises from the neighborhood aggregation operation of GNNs, of which the rationale is to iteratively refine the node embeddings with similar neighbors. Pseudo-training effect implies that replaying the TE of one node can also strengthen the prediction for its neighbors within the same computation ego-subgraph and alleviate the forgetting problem on them. More detailed discusion on pseudo-training effect is provided in Appendix A.3 and Appendix A.4. The above analysis suggests that TEs with larger computation ego-subgraphs covering more nodes may be more effective, motivating our coverage maximization sampling strategy in the next subsection, which is also empirically justified in Section 4.3.
Algorithm 1 Coverage maximization sampling
1:Input: $\mathcal{G}_{\tau}$, $\mathbb{V}_{\tau}$, $\hat{\mathbf{A}}_{\tau}$, $\pi(\cdot,\cdot;\cdot)$, sample size $n$.
2:
3:Output: Selected nodes $\mathcal{S}$
4:Initialize $\mathcal{S}=\{\}$.
5:for each $v\in\mathbb{V}_{\tau}$ do
6:     $R_{c}(\{v\})=\frac{|\{w|w\in\mathcal{G}_{\tau,v}^{sub}\}|}{|\mathbb{V}_{\tau}|}$
7:end for each
8:for each $v\in\mathbb{V}_{\tau}$ do
9:     $p_{v}$ = $\frac{R_{c}(\{v\})}{\sum_{w\in\mathbb{V}_{\tau}}R_{c}(\{w\})}$
10:end for each
11:while $n>0$ do
12:     Sample one node $v$ from $\mathbb{V}_{\tau}$ according to $\{p_{w}\mid w\in\mathbb{V}_{\tau}\}$.
13:     $\mathcal{S}=\mathcal{S}\cup\{v\}$
14:     $\mathbb{V}_{\tau}=\mathbb{V}_{\tau}\backslash\{v\}$$\triangleright$ Sampling without replacement
15:     $n\leftarrow n-1$
16:end while
3.6. Coverage Maximization Sampling
Following the above subsection, TEs with larger computation ego-subgraphs are preferred to be stored. To quantify the size of the computation ego-subgraphs, we formally define the coverage ratio of the selected TEs as the nodes covered by their computation ego-subgraphs versus the total nodes in the graph (Figure 2). Since a TE uniquely corresponds to a node, we may use ‘node’ and ‘TE’ interchangeably.
Definition 0 ().
Given a graph $\mathcal{G}$, node set $\mathbb{V}$, and function $\pi(\cdot,\cdot;\hat{\mathbf{A}})$, the coverage ratio of a set of nodes $\mathbb{V}_{s}$ is:
(9) $R_{c}(\mathbb{V}_{s})=\frac{|\cup_{v\in\mathbb{V}_{s}}\{w|w\in\mathcal{G}_{v}^{sub}\}|}{|\mathbb{V}|},$
i.e., the ratio of nodes of the entire (training) graph covered by the computation ego-subgraphs of the selected nodes (TEs).
To maximize $R_{c}(\mathcal{TEM})$, a naive approach is to start from selecting the TE with the largest coverage ratio, and then iteratively incorporate TE that increases $R_{c}(\mathcal{TEM})$ the most. However, this requires computing $R_{c}(\mathcal{TEM})$ for all candidate TEs at each iteration, which is time consuming especially on large graphs. Besides, certain randomness is also desired for the diversity of $\mathcal{TEM}$. Therefore, we propose to sample TEs based on their coverage ratio. Specifically, in task $\tau$, the probability of sampling node $v\in\mathbb{V}_{\tau}$ is $p_{v}=\frac{R_{c}(\{v\})}{\sum_{w\in\mathbb{V}_{\tau}}R_{c}(\{w\})}$. Then the nodes in $\mathbb{V}_{\tau}$ are sampled according to $\{p_{v}\mid v\in\mathbb{V}_{\tau}\}$ without replacement, as shown in Algorithm 1. In experiments, we demonstrate the correlation between the coverage ratio and the performance, which verifies the benefits revealed in Section 3.5
4. Experiments
In this section, we aim to answer the following research questions: RQ1: Whether PDGNNs-TEM works well with a reasonable buffer size? RQ2: Does coverage maximization sampling ensure a higher coverage ratio and better performance when the memory budget is tight? RQ3: Whether our theoretical results can be reflected in experiments? RQ4: Whether PDGNNs-TEM can outperform the state-of-the-art methods in both class-IL and task-IL scenarios? Due to the space limitations, only the most prominent results are presented in the main content, and more details are available in Appendix B. For simplicity, PDGNNs-TEM will be denoted as PDGNNs in this section.
4.1. Datasets
Following the public benchmark CGLB (Zhang et al., 2022a), we adopted four datasets, CoraFull, OGB-Arxiv, Reddit, and OGB-Products, with up to millions of nodes and 70 classes. Dataset statistics and task splittings are summarized in Table 8 of Appendix. In the paper, we show the results under the splittings with the largest number of tasks. More details of the datasets and results with additional configurations are provided in the Appendix B.5 and B.2, respectively.
Table 1. Performance &\&& coverage ratios of different sampling strategies and buffer sizes on OGB-Arxiv dataset ($\uparrow$ higher means better).
Table 2. Performance comparisons under class-IL on different datasets ($\uparrow$ higher means better).
Ratio of dataset /0.02 0.1 1.0 5.0 40.0
AA/Uniform samp. 12.0$\pm$1.1 24.1$\pm$1.7 42.2$\pm$0.3 50.4$\pm$0.4 53.3$\pm$0.4
Mean of feat. 12.6$\pm$0.1 25.3$\pm$0.3 42.8$\pm$0.3 50.4$\pm$0.7 53.3$\pm$0.2
Cov. Max. 14.9$\pm$0.8 26.8$\pm$1.8 43.7$\pm$0.5 50.5$\pm$0.4 53.4$\pm$0.1
Cov. ratio/Uniform samp. 0.1$\pm$0.1 0.3$\pm$0.0 3.5$\pm$0.9 15.9$\pm$1.1 84.8$\pm$1.5
Mean of feat. 0.2$\pm$0.4 0.6$\pm$0.3 7.1$\pm$0.6 29.6$\pm$1.7 91.1$\pm$0.1
Cov. Max. 0.5$\pm$1.1 2.9$\pm$1.8 22.5$\pm$1.6 46.3$\pm$0.6 92.8$\pm$0.0
C.L.T. CoraFull OGB-Arxiv Reddit OGB-Products
AA/$\uparrow$ AF/$\uparrow$ AA/$\uparrow$ AF $\uparrow$ AA/$\uparrow$ AF $\uparrow$ AA/$\uparrow$ AF $\uparrow$
Fine-tune 3.5$\pm$0.2 -95.2$\pm$0.3 4.9$\pm$0.0 -89.7$\pm$0.4 5.9$\pm$0.9 -97.9$\pm$1.8 7.6$\pm$0.6 -88.7$\pm$0.5
EWC (Kirkpatrick et al., 2017) 52.6$\pm$5.0 -38.5$\pm$7.3 8.5$\pm$0.6 -69.5$\pm$4.7 10.3$\pm$6.3 -33.2$\pm$14.6 23.8$\pm$2.2 -21.7$\pm$3.9
MAS (Aljundi et al., 2018) 6.5$\pm$1.0 -92.3$\pm$1.0 4.8$\pm$0.2 -72.2$\pm$2.6 9.2$\pm$7.6 -23.1$\pm$14.6 16.7$\pm$2.6 -57.0$\pm$24.8
GEM (Lopez-Paz and Ranzato, 2017) 7.4$\pm$0.1 -91.0$\pm$0.1 4.9$\pm$0.0 -89.8$\pm$0.2 5.0$\pm$0.0 -99.4$\pm$0.0 4.5$\pm$0.8 -94.7$\pm$0.2
TWP (Liu et al., 2021) 62.6$\pm$1.7 -30.6$\pm$2.8 6.7$\pm$1.2 -50.6$\pm$6.9 8.0$\pm$2.9 -18.8$\pm$5.1 14.1$\pm$2.1 -11.4$\pm$1.3
LwF (Li and Hoiem, 2017) 33.4$\pm$0.9 -59.6$\pm$1.2 9.9$\pm$6.7 -43.6$\pm$7.5 86.6$\pm$0.8 -9.2$\pm$0.9 48.2$\pm$0.8 -18.6$\pm$0.9
ER-GNN (Zhou and Cao, 2021) 2.9$\pm$0.0 -94.6$\pm$0.1 12.3$\pm$3.1 -79.9$\pm$3.3 20.4$\pm$2.6 -82.7$\pm$2.9 56.7$\pm$0.3 -33.3$\pm$0.5
SSM (Zhang et al., 2022c) 75.4$\pm$0.1 -9.7$\pm$0.0 48.3$\pm$0.5 -10.7$\pm$0.3 94.4$\pm$0.0 -1.3$\pm$0.0 63.3$\pm$0.1 -9.6$\pm$0.3
Joint 80.8$\pm$0.1 -3.1$\pm$0.2 56.8$\pm$0.0 -8.6$\pm$0.0 97.1$\pm$0.1 -0.7$\pm$0.1 71.5$\pm$0.1 -5.8$\pm$0.2
PDGNNs 81.9$\pm$0.1 -3.9$\pm$0.1 53.2$\pm$0.2 -14.7$\pm$0.2 96.6$\pm$0.0 -2.6$\pm$0.1 73.9$\pm$0.1 -10.9$\pm$0.2
Table 2. Performance comparisons under class-IL on different datasets ($\uparrow$ higher means better).